OpenStudy (anonymous):

Rationalize the denominator. \[10-\sqrt{3} / 10+\sqrt{3}\]

OpenStudy (anonymous):

multiply numerator and denominator by conjugate of denominator

OpenStudy (anonymous):

\[\frac{10-\sqrt{3}}{10+\sqrt{3}}\times \frac{10+\sqrt{3}}{10+\sqrt{3}}\]

OpenStudy (anonymous):

oops all that and i wrote ti wrong

OpenStudy (anonymous):

(10-root3)^2/91

OpenStudy (anonymous):

\[\frac{10-\sqrt{3}}{10+\sqrt{3}}\times \frac{10-\sqrt{3}}{10-\sqrt{3}}\]

OpenStudy (anonymous):

get what mr. knowledgeable said. on account of \[(10+\sqrt{3})(10-\sqrt{3})=10^2-\sqrt{3}^2=100-9=91\]

OpenStudy (anonymous):

now \[(10-\sqrt{3})(10-\sqrt{3})=10^2-20\sqrt{3}+3=103-20\sqrt{3}\] if you want that in the numerator